# Non-Abelian anyons in a periodically-driven Abelian model

### v2

@@ Abstract
We show that non-Abelian anyons can emerge from an Abelian topologically-ordered model subject to local time-periodic driving, with the specific example of Ising anyons in a driven toric-code model. The quasiparticle excitations of the toric code, realizing an Abelian-anyon theory, include fermionic and bosonic quasiparticles which see each other as $\pi$ fluxes, namely they result in the accumulation of a $\pi$ phase if wound around each other. The emergence of non-Abelian behaviour is explained in the fact that the modulation can engineer a band topology for the fermionic quasiparticles, by Floquet engineering complex phases of fermion pairing in the high-frequency regime. This induces the fractionalization of the fermions into Floquet-Majorana modes bound to the bosons. The latter then develop non-Abelian anyon character akin to vortices in a topological superconductor. Our findings shed light on the nonequilibrium physics of driven topologically-ordered quantum matter and may facilitate the observation of non-Abelian behaviour in engineered quantum systems.
@@
Quasiparticles with anyon quantum statistics are predicted to appear as gapped excitations in topologically-ordered quantum phases of matter, such as fractional quantum Hall and spin liquids [1, 2]. Braiding anyons of Abelian type can change the many-body wavefunction by a phase factor. Braiding non-Abelian anyons, instead, can induce a unitary transformation in a topologically-protected degenerate ground state, enabling fault-tolerant quantum computation [3, 1]. Although realizing and controlling non-Abelian topological order remains elusive, Abelian anyons have been recently observed in engineered quantum systems, ranging from ultracold atoms in optical lattices [4, 5] to Rydberg-atom arrays [6] and superconducting circuits [7]. This progress motivates the study of topologically-ordered phases out of equilibrium, such as subject to external control, which may both provide a stepping stone towards non-Abelian physics [8, 9, 10, 11, 12] and unveil novel effects, with examples given by radical chiral Floquet phases [13, 14] and fractionalized prethermalization [15].
We show that time-periodic driving can effectively turn an Abelian topologically-ordered model into a non-Abelian one. This is demonstrated with the example of Kitaev’s toric code [3], the paradigmatic $\mathds{Z}_{2}$ quantum spin liquid [2]. Abelian anyons in the toric code emerge as bosonic and fermionic quasiparticle excitations which see each other as $\pi$ fluxes: Winding a fermion around a boson (and vice versa) results, unusually, in the accumulation of a $\pi$ Peierls-like phase. The Floquet modulation of the physical spins fundamentally alters this picture by driving one Abelian-anyon type, the fermion, into a phase characterized by a non-trivial band topology. This is achieved by Floquet-engineering fermion pairing with a complex phase, mimicking key features of topological superconductors [16, 1]. Such processes induce the fractionalization of the fermionic quasiparticles into pairs of Majorana modes binding to the other (undriven) anyon type, namely the bosonic quasiparticles. We verify that these Majorana-carrying bosons behave like non-Abelian Ising anyons, known in other contexts such as $p+ip$ superconductors [16], Kitaev’s honeycomb model [8] and the Pfaffian quantum Hall state [17]. Our analysis involves analytical high-frequency expansions and the numerical determination of quasienergy spectra and of the non-Abelian exchange phases. These results shed light on the properties of nonequilibrium interacting topological phases and establish Floquet driving as a potential link between them. Moreover, they pave the way towards the development of novel protocols to Floquet-engineer non-Abelian anyons [18, 19] building on the control of Abelian phases recently attained [6] or predicted [20, 21, 22].
@@
The driven model. The Abelian toric-code model [3, 23] describes two-level systems (spins) on the bonds of a square $L\times L$ lattice [Fig. 1(a)], with $L$ even hereafter, and ruled by the Hamiltonian
@@
The labels $p$ and $v$ denote lattice plaquettes and vertices, $g$ is a coupling constant, $\hat{A}_{v}=\prod_{i\in v}\hat{X}_{i}$ and $\hat{B}_{p}=\prod_{i\in p}\hat{Z}_{i}$ are four-spin interactions and $\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}$ are Pauli matrices for the $i$th spin. When convenient, we will also interpret $p=(p_{x},p_{y})$ and $v=(v_{x},v_{y})$ as coordinates of two separate plaquette and vertex lattices, respectively, both with lattice vectors $x=(1,0)$ and $y=(0,1)$. All four-spin operators composing $\hat{H}_{\rm tc}$ commute with each other, $[\hat{A}_{v},\hat{A}_{v^{\prime}}]=[\hat{A}_{v},\hat{B}_{p}]=[\hat{B}_{p},\hat{B}_{p^{\prime}}]=0$, making the model exactly solvable. Since $\hat{B}^{2}_{p}=\hat{A}^{2}_{v}=\mathds{1}$, the eigenstates can be labelled by the eigenvalues $A_{v}=\pm 1$ and $B_{p}=\pm 1$ for each vertex and plaquette. On a torus, where it can be shown that $\prod_{v}\hat{A}_{v}=\prod_{p}\hat{B}_{p}=\mathds{1}$, the eigenstates are further divided into four superselection sectors, identified by the ±$\pm 1$ eigenvalues of two operators $\hat{O}_{x}=\prod_{i\in\Lambda_{x}}{\hat{X}_{i}}$ and $\hat{O}_{y}=\prod_{i\in\Lambda_{y}}{\hat{X}_{i}}$ spanning the torus along horizontal ($\Lambda_{x}$) and vertical ($\Lambda_{y}$) noncontractible loops and commuting with all $\hat{A}_{v}$ and $\hat{B}_{p}$ [Fig. 1(a)]. The model has four degenerate ground states, satisfying $A_{v}=B_{p}=1$, which are distinguished by the eigenvalues of $\hat{O}_{x}$ and $\hat{O}_{y}$. Flipping one eigenvalue, $A_{v}=-1$ ($B_{p}=-1$), costs an excitation energy $g$ and is interpreted as the creation of a pointlike quasiparticle $e$ ($m$) on top of the ground state $\mathds{1}_{\rm tc}$, which occupies the corresponding vertex $v$ (plaquette $p$). While $e$ and $m$ are hardcore bosons under self-exchange, they are mutual semions: the wavefunction picks up a Peierls-like $\pi$ phase when $m$ is wound around $e$, or vice versa. A bound state $\psi=e\times m$ of an $e$ and $m$ particle, costing an energy $\mu_{\psi}=2g$, is then a fermion, while still being a semion with respect to a separate $e$ or $m$. These quasiparticles, including $\mathds{1}_{\rm tc}$, represent the four anyon types in the toric code. They are Abelian, since their non-trivial fusion has a single possible outcome: $e\times m=\psi$, $e\times\psi=m$, $m\times\psi=e.$
We add to $\hat{H}_{\rm tc}$ a drive in order to set the $\psi$-fermions in motion and imbue them with a topological band structure. Non-Abelian character will emerge from the key interplay between the fermion band topology, achieved through Floquet engineering, and the semion relation between the $\psi$ and $e$ Abelian anyons, provided by the underlying toric code. The drive is chosen as
@@
The subscripts $({p,j})$, with $j\in\{B,R,T,L\}$ denote the spins at the bottom, right, top or left of $p$, respectively [Fig. 1(a)]. Two-body couplings of the form $\hat{X}_{i}\hat{Z}_{j}$ on adjacent spins induce $\psi$ tunnelling and pairing, as sketched in Fig. 1(a): they displace a joint pair of neighbouring $e$ and $m$ composing $\psi$, as well as create and annihilate pairs of them. Indeed, since $\hat{X}_{i}$ anticommutes with $\hat{B}_{p}$ on neighbouring plaquettes $p$ and $p^{\prime}$ while commuting with $\hat{A}_{v}$, its action on an eigenstate flips the eigenvalues $B_{p}$ and $B_{p^{\prime}}$, while preserving the joint $m$-anyon parity $(-1)^{B_{p}+B_{p^{\prime}}}$. Hence, if a $m$-particle composing a fermion is present at $p$, it tunnels to $p^{\prime}$ (and vice versa). If no fermion or a pair is present, the associated pair of $m$ will be created or annihilated. An equivalent analysis applies to $e$-motion on nearby vertices induced by $\hat{Z}_{j}$. The control functions $d_{x}(t)$ and $d_{y}(t)$ govern processes along the horizontal and vertical direction of the lattice, respectively.
@@
Figure 1: (a) Toric-code spin lattice. The $m$ anyons live on plaquettes $p$ (orange), while $e$ anyons live on vertices $v$ (blue), and their bound state forms a fermion $\psi$. The eigenvalues of $\hat{O}_{x}$ and $\hat{O}_{y}$ define four superselection sectors in the undriven system. The operators $\hat{X}_{i}\hat{Z}_{j}$ depicted induce fermion tunnelling and pairing. Dashed lines indicate periodic boundaries. (b) Lattice for $\psi$ fermions in the quasiparticle mapping. The controls $d_{x}$ and $d_{y}$ drive fermion tunnelling and pairing. Vertical processes include a boson operator $\hat{\mathcal{S}}_{p}^{(y)}$ reproducing the mutual statistics of $\psi$ and $e$. (c) Drive-assisted fermion pairing, restoring otherwise off-resonant ($J\ll\mu_{\psi}$) processes with rate $\Delta$ and phase $\phi_{r}$.
@@
Quasiparticle picture. To discuss the impact of the drive on the $\psi$ quasiparticles, we adopt the quasiparticle picture of Ref. [24] (details are given in the Supplemental Material (SM) [25]). The spins’ Hilbert space is mapped to the tensor product of Fock spaces for $e$ hardcore-boson and $\psi$ fermion occupations, and of a four-dimensional Hilbert space reflecting four superselection sectors. The spin operators map to creation and annihilation operators of bosons on vertices ($\hat{b}_{v}^{\dagger},\hat{b}_{v}$) and of fermions on plaquettes ($\hat{f}_{p}^{\dagger},\hat{f}_{p}$), with the toric-code ground state |$\ket{0_{\rm tc}}$ representing the quasiparticle vacuum. The Hamiltonian $\hat{H}_{\rm tc}$ maps to $\hat{H}_{e}+\hat{H}_{e\psi}^{(0)}$ with $\hat{H}_{e}=g\sum_{v}\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and $\hat{H}_{e\psi}^{(0)}=2g\sum_{p}[1-\hat{b}^{\dagger}_{v(p)}\hat{b}_{v(p)}]\hat{f}_{p}^{\dagger}\hat{f}_{p}$, describing quasiparticles at rest [$v(p)$ denotes the vertex to the bottom-left of $p$]. The drive $\hat{H}_{d}(t)$ maps to $\hat{H}_{e\psi}^{(d)}(t)=-\sum_{p,r\in\{x,y\}}d_{r}(t)\hat{\mathcal{S}}_{p}^{(r)}(\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.})$, describing fermion tunnelling and pairing but also containing a coupling to the $e$-bosons through $\hat{\mathcal{S}}_{p}^{(x)}=1$ and $\hat{\mathcal{S}}_{p}^{(y)}=\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}$. Here, $\mathrm{R}[v(p)]$ denotes the vertices to the right of the spin shared by the plaquettes $p$ and $p+y$, see Fig. 1(b). This coupling explicitly reproduces the accumulation of Peierls-like phase as the anyons move with respect to each other: Considering a fermion tunnelling on a loop encircling a $e$ particle, $\hat{\mathcal{S}}_{p}^{(r)}$ indeed yields the accumulation of $\pi$ phase. Since the drive does not induce boson motion, all bosonic terms are diagonal in the Fock basis |$\ket{\vec{n}^{e}}=(\hat{b}_{v_{1}}^{\dagger})^{n_{1}}\dots(\hat{b}_{v_{N}}^{\dagger})^{n_{N}}\ket{0_{\rm tc}}$ describing $e$ occupations. We can then analyze the fermion Hamiltonian $\hat{H}_{\psi}(t)=\bra{\vec{n}^{e}}\hat{H}_{e\psi}^{(0)}+\hat{H}_{e\psi}^{(d)}(t)\ket{\vec{n}^{e}}$ corresponding to a fixed distribution of $e$ particles. The presence of $e$-bosons at vertices $\{v\}$ is assumed to derive from preparing the ground state of a slightly modified Hamiltonian (1) with inverted coupling $g\to\tilde{g}=-g$ for the vertices $\{v\}$. The Hamiltonian $\hat{H}_{\psi}(t)$ then reads
@@
It describes spinless superconducting fermions coupled, through $\hat{\mathcal{S}}_{p}^{(r)}$, to an effective flux given by a background distribution of $e$ particles (also interpretable as $\mathbb{Z}_{2}$ vortices of a deconfined $\mathbb{Z}_{2}$ gauge field [2, 26]).
Floquet engineering. The above ingredients suggest a potential analogy to topological superconductors [16, 1, 27]. In such systems, non-Abelian anyons are predicted to emerge as vortices ($\sigma$) carrying a Majorana zero-mode [28]. Majorana modes attached to topological defects result from the fractionalization of the superconducting fermions when their bulk bands are gapped and topological, as is the case for pairing in so-called $p+ip$ symmetry [1]. In a lattice model as in (3), this corresponds to a complex-valued pairing coupling $\Delta_{r}f_{p}^{\dagger}f_{p+r}^{\dagger}$ along $x$ and $y$ of the form $(\Delta_{x},\Delta_{y})=(\Delta,i\Delta)$ [26]. The resulting so-called Ising anyons $\sigma$, $\mathds{1}_{\rm is}$ and $\psi$ feature non-trivial fusion rules $\sigma\times\sigma=\mathds{1}_{\rm is}+\psi$, $\psi\times\psi=\mathds{1}_{\rm is}$, $\psi\times\sigma=\sigma$ [1, 8]. The multiple fusion channels for the vortex $\sigma$, yielding either the vacuum $\mathds{1}_{\rm is}$ or a fermion $\psi$, qualify them as non-Abelian. We will show that the driven toric code reproduces faithfully this physics and non-Abelian anyons, with $e$ particles behaving like $\sigma$ vortices, for appropriate driving functions that Floquet-engineer $p+ip$ fermion pairing.
Since the control can only address the physical spins, rather than the quasiparticle pairing terms directly, no choice of time-independent $d_{x}$ and $d_{y}$ in Eq. (2) can achieve the goal. Indeed, the functions $d_{r}(t)$ need be real-valued and they induce tunnelling and pairing with equal real rate. We overcome these limitations employing a periodic modulation $d_{r}(t)=d_{r}(t+T)$ and Floquet engineering [29, 30]. The driving functions are chosen as
@@
[…]
$\hat{H}_{F}=$ $-\sum_{p}\Big{[}\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}$
@@
At time $T$, the transformation $\hat{R}(t)$ reduces to the total fermion parity, $\hat{R}(T)=(-1)^{\sum_{p}\hat{f}^{\dagger}_{p}\hat{f}_{p}}$, which is conserved by the driven (and thus also the Floquet) Hamiltonian. Since quasiparticles in the toric code can only be created in pairs, only even-parity states are physical. Given that $R(T)=1$ in their subspace, their stroboscopic dynamics is ruled by $\hat{H}_{F}$ of Eq. (5). We obtain the desired $p+ip$ pairing in $\hat{H}_{F}$ by choosing a circular-like shaking, with phase delay $\phi_{y}-\phi_{x}=\pi/2$ between horizontal and vertical modulations. The detuning of the drive from the excitation energy of a fermion pair in the toric code controls the effective chemical potential $\mu$, and the fermions are coupled to the $e$ particles via ${\mathcal{S}}_{p}^{(r)}$.
@@
Figure 2: (a) Splitting between fermion ground states for $\mu=-2J$ in different superselection sectors, having energy $E_{\rm GS}^{(x,y)}$ [$(x,y)$ denotes the boundary conditions, either periodic ($P$) or antiperiodic ($A$) along $x$ and $y$] in the high-frequency regime ($J=\Delta=0.1\mu_{\psi}$, solid lines and circles) and in the anomalous phase ($J=\Delta=0.15\mu_{\psi}$, dashed lines and squares). (b) Floquet-BdG quasienergies on a $L_{x}\times L_{y}=20\times 200$ cylinder, for $J=\Delta=0.1\mu_{\psi}$, and (c) on a $(A,A)$-torus in the presence of a $e$ pair, for increasing driving amplitude ($\mu=-2J$, $L=24$). The color-scale indicates the inverse participation ratio of the corresponding eigenstate. (d) Paths $P$ and $P^{\prime}$ involved in the vortex-exchange protocol. (e) Vortex exchange phase $\theta_{P}-\theta_{P^{\prime}}$ in the $\mathds{1}_{\rm is}$ (bottom) and $\psi$ (top) sectors at $\omega=2(\mu_{\psi}-2J)$, for different driving amplitude and system size.

### v3

# Non-Abelian Anyons in Periodically Driven Abelian Spin Liquids

@@ Abstract
We show that non-Abelian anyons can emerge from an Abelian topologically-ordered system subject to local time-periodic driving. This is illustrated with the toric-code model, as the canonical representative of a broad class of Abelian topological spin liquids. The Abelian anyons in the toric code include fermionic and bosonic quasiparticle excitations which see each other as $\pi$ fluxes, namely they result in the accumulation of a $\pi$ phase if wound around each other. Non-Abelian behaviour emerges because the Floquet modulation can engineer a non-trivial band topology for the fermions, inducing their fractionalization into Floquet-Majorana modes bound to the bosons. The latter then develop non-Abelian character akin to vortices in topological superconductors, realizing Ising topological order. Our findings shed light on the nonequilibrium physics of driven topologically-ordered quantum matter and may facilitate the observation of non-Abelian behaviour in engineered quantum systems.
@@
Quasiparticles with anyon quantum statistics are predicted to appear as gapped excitations in topologically-ordered (TO) quantum phases of matter, such as fractional quantum Hall and spin liquids [1, 2]. Braiding anyons of Abelian type can change the many-body wavefunction by a phase factor. Braiding non-Abelian anyons, instead, can induce a unitary transformation in a topologically-protected degenerate ground state, enabling fault-tolerant quantum computation [3, 4, 1]. Abelian anyons have been recently observed and manipulated in various engineered quantum systems, ranging from ultracold atoms in optical lattices [5, 6] to Rydberg-atom arrays [7] and superconducting circuits [8, 9], and first realizations of non-Abelian TO states have been reported in trapped ions quantum processors [10]. This progress motivates the study of TO phases out of equilibrium, such as subject to external control, which may both provide a stepping stone towards non-Abelian physics [4, 11, 12, 13, 9] and unveil novel effects, with examples being radical chiral Floquet phases [14, 15] and fractionalized prethermalization [16].
Here we show that time-periodic driving can effectively turn an Abelian TO system into a non-Abelian one. We use as a testbed the paradigmatic toric-code model [3], which is believed to capture the fundamental low-energy physics of a large class of Abelian topological spin liquids [2, 17] while remarkably being exactly soluble. This argues for the broad applicability of our results while enabling a clear illustration of the desired effect through exact methods. Abelian anyons in $\mathds{Z}_{2}$ spin liquids like the toric code emerge as bosonic and fermionic quasiparticle excitations which see each other as $\pi$ fluxes: Winding a fermion around a boson (and vice versa) results, unusually, in the accumulation of a $\pi$ phase. We show that a local Floquet modulation can alter the picture by driving one anyon type, the fermion, into a superconducting phase with a non-trivial band topology. This induces the fractionalization of the fermions into pairs of Majorana modes which bind to the other anyon type (the boson), realizing non-Abelian Ising TO [1, 4, 18]. This is predicted through analytical high-frequency expansions and verified by numerical computation of the non-Abelian exchange phases. Our results shed light on the properties of driven interacting topological systems and promote the development of protocols to Floquet-engineer non-Abelian anyons [19, 20] from Abelian phases, further motivated by the recent experimental realizations [7, 8, 21] and theoretical proposals [22, 23, 24, 25] of toric-code TO in quantum simulators. Complementary to achievements in the manipulation of individual TO states [8, 9, 10], a Floquet-engineering approach has the appeal of realizing the background Hamiltonian, which stabilizes the desired TO states, makes anyons well-defined quasiparticles [26] and enables the study of dynamics.
@@
The driven model.— The toric-code model [3, 27] describes spin-1/2 systems on the bonds of a square $L\times L$ lattice [Fig. 1(a)] ($L$ even hereafter) with Hamiltonian
@@
The labels $p$ and $v$ denote lattice plaquettes and vertices, $g$ is a coupling constant, $\hat{A}_{v}=\prod_{i\in v}\hat{X}_{i}$ and $\hat{B}_{p}=\prod_{i\in p}\hat{Z}_{i}$ are four-spin interactions and $\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}$ are Pauli matrices for the $i$th spin. When convenient, we will also interpret $p=(p_{x},p_{y})$ and $v=(v_{x},v_{y})$ as coordinates of two separate plaquette and vertex lattices, respectively, both with lattice vectors $x=(1,0)$ and $y=(0,1)$. All four-spin operators in (1) commute with each other, making the model exactly solvable. The eigenstates can be labelled by the eigenvalues $A_{v}=\pm 1$ and $B_{p}=\pm 1$ for each vertex and plaquette. On a torus, since $\prod_{v}\hat{A}_{v}=\prod_{p}\hat{B}_{p}=\mathds{1}$, a complete set of observables is obtained by including two operators $\hat{O}_{x}=\prod_{i\in\Lambda_{x}}{\hat{X}_{i}}$ and $\hat{O}_{y}=\prod_{i\in\Lambda_{y}}{\hat{X}_{i}}$ spanning noncontractible loops [$\Lambda_{x}$ and $\Lambda_{y}$ in Fig. 1(a)] and commuting with all $\hat{A}_{v}$ and $\hat{B}_{p}$. The ±$\pm 1$ eigenvalues of $\hat{O}_{x/y}$ define four superselection sectors, leading to four degenerate ground states satisfying $A_{v}=B_{p}=1$. Flipping one eigenvalue, $A_{v}=-1$ ($B_{p}=-1$), costs an excitation energy $g$ and is interpreted as the creation of a pointlike quasiparticle $e$ ($m$) on top of the ground state $\mathds{1}_{\rm tc}$, which occupies the corresponding vertex $v$ (plaquette $p$). While $e$ and $m$ are hardcore bosons under self-exchange, they are mutual semions: the wavefunction picks up a Peierls-like $\pi$ phase when $m$ is wound around $e$, or vice versa. A bound state $\psi=e\times m$, costing an energy $\mu_{\psi}=2g$, is then a fermion, while still being a semion with respect to a separate $e$ or $m$. These quasiparticles, including $\mathds{1}_{\rm tc}$, represent the four anyon types in the toric code. They are Abelian, since their non-trivial fusion has a single possible outcome: $e\times m=\psi$, $e\times\psi=m$, $m\times\psi=e.$
We add a drive to $\hat{H}_{\rm tc}$ to induce $\psi$-fermion motion and realize a topological band structure. Non-Abelian character will emerge from the interplay between the fermion band topology, achieved through Floquet engineering, and the semion relation between the $\psi$ and $e$ Abelian anyons, provided by the underlying toric code. The drive addresses neighbouring spins as
@@
The subscripts $({p,j})$, with $j\in\{B,R,T,L\}$ denote the spins at the bottom, right, top or left of $p$, respectively [Fig. 1(a)]. We expect drives of even simpler structure to achieve the desired effect, but (2) has the key merit that, while still being local, it allows for a transparent analytical treatment and efficient numerical simulation of the time-dependent system, as we shall see, thus enabling a clear illustration of the target physics. This is an important advantage: investigating the toric code with even static magnetic-field perturbations generically requires sophisticated methods not directly applicable to the driven case [28, 29, 30]. Two-spin couplings of the form $\hat{X}_{i}\hat{Z}_{j}$ on adjacent spins induce $\psi$ tunnelling and pairing: they displace a pair of neighbouring $e$ and $m$ composing $\psi$, as well as create and annihilate pairs of them. Indeed, since $\hat{X}_{i}$ anticommutes with $\hat{B}_{p}$ on neighbouring plaquettes $p$ and $p^{\prime}$ while commuting with $\hat{A}_{v}$, its action on an eigenstate flips the eigenvalues $B_{p}$ and $B_{p^{\prime}}$, while preserving the joint $m$-anyon parity $(-1)^{B_{p}+B_{p^{\prime}}}$. Hence, if a $m$-particle composing a fermion is present at $p$, it tunnels to $p^{\prime}$ (and vice versa). If no fermion or a pair is present, the associated pair of $m$ will be created or annihilated. An equivalent analysis applies to $e$-motion on nearby vertices induced by $\hat{Z}_{j}$. The functions $d_{x}(t)$ and $d_{y}(t)$ control processes along the horizontal and vertical direction, respectively.
@@
Figure 1: (a) Toric-code spin lattice. The $m$ anyons live on plaquettes $p$ (orange), while $e$ anyons live on vertices $v$ (blue), and their bound state forms a fermion $\psi$. The operators $\hat{X}_{i}\hat{Z}_{j}$ depicted induce fermion tunnelling and pairing. Dashed lines indicate periodic boundaries. (b) Corresponding lattice for $\psi$ fermions in the quasiparticle mapping, where $d_{x}$ and $d_{y}$ control fermion tunnelling and pairing. Vertical processes include a boson operator $\hat{\mathcal{S}}_{p}^{(y)}$ reproducing the mutual statistics of $\psi$ and $e$. (c) Drive-assisted fermion pairing, restoring otherwise off-resonant ($J\ll\mu_{\psi}$) processes with rate $\Delta$ and phase $\phi_{r}$.
@@
Quasiparticle picture.— With the chosen form of the modulation, the driven model can be mapped exactly to a problem of driven noninteracting $\psi$ fermions coupled to static $e$ bosons, which allows us to focus on the impact of the drive on the $\psi$ anyons. To this end, we adopt the quasiparticle picture of Ref. [31] (details are given in the Supplemental Material (SM) [32]). The spins’ Hilbert space is mapped to the tensor product of Fock spaces for $e$ hardcore-boson and $\psi$ fermion occupations, and of a four-dimensional Hilbert space reflecting four superselection sectors. The spin operators map to creation and annihilation operators of bosons on vertices ($\hat{b}_{v}^{\dagger},\hat{b}_{v}$) and of fermions on plaquettes ($\hat{f}_{p}^{\dagger},\hat{f}_{p}$), with the toric-code ground state |$\ket{0_{\rm tc}}$ representing the quasiparticle vacuum. The Hamiltonian $\hat{H}_{\rm tc}$ maps to $\hat{H}_{e}+\hat{H}_{e\psi}^{(0)}$ with $\hat{H}_{e}=g\sum_{v}\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and $\hat{H}_{e\psi}^{(0)}=2g\sum_{p}[1-\hat{b}^{\dagger}_{v(p)}\hat{b}_{v(p)}]\hat{f}_{p}^{\dagger}\hat{f}_{p}$, describing quasiparticles at rest [$v(p)$ denotes the vertex to the bottom-left of $p$]. The drive $\hat{H}_{d}(t)$ maps to $\hat{H}_{e\psi}^{(d)}(t)=-\sum_{p,r\in\{x,y\}}d_{r}(t)\hat{\mathcal{S}}_{p}^{(r)}(\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.})$, describing fermion tunnelling and pairing but also containing a coupling to the $e$-bosons through $\hat{\mathcal{S}}_{p}^{(x)}=1$ and $\hat{\mathcal{S}}_{p}^{(y)}=\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}$. Here, $\mathrm{R}[v(p)]$ denotes the vertices to the right of the spin shared by the plaquettes $p$ and $p+y$, see Fig. 1(b). This coupling explicitly reproduces the semion mutual statistics: $\hat{\mathcal{S}}_{p}^{(r)}$ yields the accumulation of $\pi$ phase when a fermion encircles a $e$ particle. Since the drive does not induce boson motion, all bosonic terms are diagonal in the Fock basis |$\ket{\vec{n}^{e}}=(\hat{b}_{v_{1}}^{\dagger})^{n_{1}}\dots(\hat{b}_{v_{N}}^{\dagger})^{n_{N}}\ket{0_{\rm tc}}$ describing $e$ occupations. We can then analyze the fermion Hamiltonian $\hat{H}_{\psi}(t)=\bra{\vec{n}^{e}}\hat{H}_{e\psi}^{(0)}+\hat{H}_{e\psi}^{(d)}(t)\ket{\vec{n}^{e}}$ corresponding to a fixed distribution of $e$ particles. The potential presence of $e$-bosons at vertices $\{v\}$ is assumed to derive from preparing the ground state of a slightly modified Hamiltonian (1) with inverted coupling $g\to\tilde{g}=-g$ for those vertices. The Hamiltonian $\hat{H}_{\psi}(t)$ then reads
@@
where $\mathcal{S}_{p}^{(r)}=\bra{\vec{n}^{e}}\hat{\mathcal{S}}_{p}^{(r)}\ket{\vec{n}^{e}}$. It describes spinless superconducting fermions coupled, through $\hat{\mathcal{S}}_{p}^{(r)}$, to an effective gauge flux given by a background distribution of $e$ particles. For a fixed $e$-boson distribution, the driven toric code then maps to a problem of driven noninteracting fermions.
Floquet engineering.— The above ingredients suggest a potential analogy to topological superconductors [33, 1, 34], where non-Abelian anyons emerge as vortices ($\sigma$) carrying Majorana zero modes [35]. The latter result from the fractionalization of the fermions occurring when their bulk bands are gapped and topological, as is the case for pairing in so-called $p+ip$ symmetry [1]. In the lattice model (3), this corresponds to a complex-valued pairing coupling $\Delta_{r}f_{p}^{\dagger}f_{p+r}^{\dagger}$ with $(\Delta_{x},\Delta_{y})=(\Delta,i\Delta)$ [17]. These so-called Ising anyons $\sigma$, $\mathds{1}_{\rm is}$ and $\psi$ feature non-trivial fusion rules $\sigma\times\sigma=\mathds{1}_{\rm is}+\psi$, $\psi\times\psi=\mathds{1}_{\rm is}$, $\psi\times\sigma=\sigma$ [1, 4]. The multiple fusion channels for the vortex $\sigma$, yielding either the vacuum $\mathds{1}_{\rm is}$ or a fermion $\psi$, qualify them as non-Abelian. We will show that the driven toric code reproduces faithfully this physics, with $e$ particles behaving like $\sigma$ vortices, for appropriate driving functions that Floquet-engineer complex fermion pairing.
Since the drive controls the physical spins, rather than the quasiparticle pairing terms directly, no choice of time-independent $d_{x}$ and $d_{y}$ in Eq. (2) can achieve the goal. Indeed, the functions $d_{r}(t)$ need be real-valued and they induce tunnelling and pairing with equal real rate. We overcome these limitations employing a periodic modulation $d_{r}(t)=d_{r}(t+T)$ and Floquet engineering [36, 37]. The driving functions are chosen as
@@
with frequency $\omega=2\pi/T$ and amplitudes $J$ and $\Delta$ much smaller than the fermion chemical potential, $J,\Delta\ll\mu_{\psi}$. They only differ in their phase $\phi_{r}$. Consider first the limit $\Delta=0$ in Eq. (4). Since the energy $2\mu_{\psi}$ required for pair creation and annihilation is much larger than $J$, pairing processes with a real coupling $J$ in (3) are far off-resonant and effectively suppressed, leaving only quasiparticle tunnelling. They are restored via the modulation with $\Delta\neq 0$ and a frequency $\omega$ quasiresonant with $2\mu_{\psi}$ [Fig. 1(c)]. The advantage is that a complex phase can be attributed to the ‘photon-assisted’ pairing coupling, depending on the phases $\phi_{r}$ of the drives. We choose $\omega=2(\mu_{\psi}+\mu)$, allowing for a small detuning $2\mu$. Applying Floquet theory [38, 39], the stroboscopic dynamics induced by the time-periodic Hamiltonian (3) in steps of $T$ is captured as $\hat{U}(nT)=e^{-i\hat{H}_{F}nT}$ by a time-independent Floquet Hamiltonian $\hat{H}_{F}$, which can be approximated from $\hat{H}_{\psi}(t)$ through high-frequency expansions, in the regime $\omega\gg J,\Delta,\mu$ considered here [40, 36, 41, 42, 37]. To leading order in $\omega^{-1}$ and in a frame defined by $\hat{R}(t)=\exp[-it(\omega/2)\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}]$ [32], $\hat{H}_{F}$ is approximated by the time-averaged Hamiltonian
$\hat{H}_{\rm avg}$ $=-\sum_{p}\Big{[}\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}$
@@
At time $T$, $\hat{R}(T)=(-1)^{\sum_{p}\hat{f}^{\dagger}_{p}\hat{f}_{p}}$ reduces to the total fermion parity, which is a conserved quantity. Since quasiparticles in the toric code can only be created in pairs, only even-parity states are physical and $\hat{R}(T)=1$ in their subspace. The desired $p+ip$ pairing in $\hat{H}_{\rm avg}$ is obtained by choosing a circular-like shaking, with phase delay $\phi_{y}-\phi_{x}=\pi/2$ between horizontal and vertical modulations. The effective chemical potential $\mu$ is controlled by the detuning of the drive from the excitation energy of a fermion pair in the toric code, and the fermions are coupled to the $e$ particles via ${\mathcal{S}}_{p}^{(r)}$.
@@
Figure 2: (a) Splitting between fermion ground states for $\mu=-2J$ in different superselection sectors, having energy $E_{\rm GS}^{(x,y)}$ [$(x,y)$ denotes the boundary conditions, either periodic ($P$) or antiperiodic ($A$) along $x$ and $y$] in the high-frequency regime ($J=\Delta=0.1\mu_{\psi}$, solid lines and circles) and in the anomalous phase ($J=\Delta=0.15\mu_{\psi}$, dashed lines and squares). (b) Floquet-BdG quasienergies on a $L_{x}\times L_{y}=20\times 200$ cylinder, for $J=\Delta=0.1\mu_{\psi}$, and (c) on a $(A,A)$-torus in the presence of a $e$ pair, for increasing driving amplitude ($\mu=-2J$, $L=24$). The color-scale indicates the inverse participation ratio of the corresponding eigenstate. (d) Paths $P$ and $P^{\prime}$ involved in the vortex-exchange protocol. (e) Vortex exchange phase $\theta_{P}-\theta_{P^{\prime}}$ in the $\mathds{1}_{\rm is}$ (bottom) and $\psi$ (top) sectors at $\omega=2(\mu_{\psi}-2J)$, for different driving amplitudes and system sizes.